From the following data, calculate the total market value of the firm. Earnings before interest = $0.1 million, D = $0.2 million, interest on debt = 10% p.a., cost of equity capital = 20%, the dividend payout ratio = 0.5, and assume no taxes
Answer #1

Earnings before interest = $0.1 million, D = $0.2 million, interest on debt = 10% p.a., cost of equity capital = 20%, the dividend payout ratio = 0.5, and assume no taxes

Value of the firm =  OFCF ​÷(k−g)

where:OFCF​=operating free cash flow

k=discount rate, in this case WACC ( weighted average cost of capital )

g=expected growth rate in OFCF​

g=RR×ROIC

where:

RR=average retention rate, or (1 - dividend payout ratio) = (1- 0.5) = 0.5 = 50%

Return on the invested capital ROIC=EBIT(1−tax)÷total capital​

Tax = 0

Total Capital not mentioned here

Equity Capital is Not mentioned here. Assuming Equity Capital same as debt = then Equity Capital = $0.2 million

ROIC = 0.1 / 0.2+0.2= 25%

g = RR X ROIC = 50% X 25% = 12.5%

K = WACC = WACC is given by the formula (Weight of Debt  x Rate of Debt ) + (Weight of Equity X Rate of Equity )

= 50% * 10%. + 50% * 20% = 15% =

so k as stated above is 15%  

Taking EBIT as Operating Cash flow ( as no other information is given ) = $ 0.1 m ( as given above )

Value of the firm =   OFCF​÷(k−g) = $ 0.1 / ( 15% - 12.5%) = $ 4 million ( answer )